# Intrinsic superconducting diode effect and nonreciprocal superconductivity in rhombohedral graphene multilayers

### v1

@@ Abstract
Rhombohedral tetralayer graphene has recently emerged as an exciting platform for a possible chiral superconducting state. Here, we theoretically demonstrate and study the emergence of nonreciprocal superconductivity and an intrinsic superconducting diode effect in this system. Our results are based on a fully self-consistent framework for determining the superconducting order parameter from a Kohn-Luttinger mechanism to superconductivity and show that large diode efficiencies, $\sim 60$%, are achievable and highly tunable by an external displacement field. Moreover, we also find that the diodicity shows a characteristic angular dependence with multiple enhanced lobes, which depend on the Fermi surface structure of the underlying normal state. Hence, our results suggest that the intrinsic superconducting diode effect could provide insights into the type of Fermi surface topology from which superconductivity arises.
@@
Non-reciprocal phenomena in superconducting systems that break time-reversal and inversion symmetry have recently attracted a great deal of attention [1]. A prominent example is the superconducting diode effect (SDE), characterized by a critical current that depends on the current-bias direction [2, 3, 4, 5, 6, 7, 8]. One established route to achieving such a SDE relies on applying an external magnetic field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or flux [19, 20, 21, 22, 23, 24, 25, 26] to break time-reversal symmetry. However, it is also possible for certain materials to break the relevant symmetries at the microscopic level, giving rise to an intrinsic SDE [27, 28, 29, 30, 31, 32, 33]. Such an intrinsic SDE could enable field-free non-reciprocal circuit elements [34, 35, 36, 37, 38, 39] and serve as a novel probe for symmetry-breaking in superconducting materials.
@@
In an exciting recent development, superconductivity has been discovered in rhombohedral tetralayer graphene (RTLG) [40, 41]. A unique aspect of RTLG is that a superconducting phase can emerge from a spin- and valley-polarized “quarter-metal” parent state and persists under large in-plane magnetic fields [40]. These observations point to a possible spin-polarized pairing with a chiral $p+ip$ superconducting order parameter, as suggested by theoretical works  [42, 43, 44, 45, 46, 47, 48, 49, 50]. Notably, both time-reversal symmetry (due to the quarter-metal normal state and chiral superconductivity) and inversion symmetry (due to an applied displacement field and trigonal warping) can be broken in RTLG. These symmetry-breaking properties make RTLG a promising platform for non-reciprocal superconducting phenomena. Indeed, an observation of an intrinsic SDE in RTLG has recently been reported [40].
@@
Figure 1: (a) Dispersion of RTLG along the $k_{x}$-direction for different displacement fields, $\Delta_{1}$, showing a band flattening around $k_{x}=0$. (b) Density of states (DOS) of RTLG as a function of $\Delta_{1}$ and the carrier density, $n_{e}$ (relative to the charge neutrality point), showing a significant enhancement due to the band flattening. (c) Fermi surfaces for ($\Delta_{1},n_{e}$), from left to right, set to $(72\,\text{meV},0.574)$, $(67.5\,\text{meV},0.386)$, $(65\,\text{meV},0.414)$, $(63\,\text{meV},0.254)$.
@@
Here, we theoretically demonstrate and study the emergence of nonreciprocal superconductivity, characterized by an asymmetric quasiparticle dispersion, and an intrinsic SDE in RTLG. Our results are based on a fully self-consistent framework for determining the superconducting order parameter from a Kohn-Luttinger mechanism to superconductivity and show that large diode efficiencies, $\sim 60$%, are achievable and highly tunable by an external displacement field. Moreover, we also find that the diodicity shows a characteristic angular dependence with multiple enhanced lobes, which depend on the Fermi surface topology of the underlying normal state. Hence, our results suggest that the intrinsic SDE could provide insights into the type of Fermi surface topology from which superconductivity arises.
@@
Model. We consider a minimal model for spin- and valley-polarized RTLG. The non-interacting part of the Hamiltonian, $H_{0}$, for describing the band structure near the $K$ or $K^{\prime}$ valley (labeled by $\tau=\pm$) is written in the basis $(A_{1},B_{1},A_{2},B_{2},A_{3},B_{3},A_{4},B_{4})$, where $A_{\ell}$ and $B_{\ell}$ denote the sublattices of the $\ell^{\text{th}}$ layer. It is given by [51],
@@
$\begin{pmatrix}D-\delta&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi&0&\frac{\gamma_{2}}{2}&0&0\\ v_{0}\pi&D&\gamma_{1}&v_{4}\pi^{\dagger}&0&0&0&0\\ v_{4}\pi&\gamma_{1}&\frac{D}{3}&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi&0&\frac{\gamma_{2}}{2}\\ v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&\frac{D}{3}&\gamma_{1}&v_{4}\pi^{\dagger}&0&0\\ 0&0&v_{4}\pi&\gamma_{1}&-\frac{D}{3}&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi\\ \frac{\gamma_{2}}{2}&0&v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&-\frac{D}{3}&\gamma_{1}&v_{4}\pi^{\dagger}\\ 0&0&0&0&v_{4}\pi&\gamma_{1}&\delta-D&v_{0}\pi^{\dagger}\\ 0&0&\frac{\gamma_{2}}{2}&0&v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&-D-\delta\end{pmatrix},$
where $\pi=\tau k_{x}+ik_{y}$ and the momentum $\boldsymbol{k}=(k_{x},k_{y})$ is measured relative to $K$ or $K^{\prime}$ valley. The $k_{x}$-direction correspond to the $\Gamma$-$K$ direction of the full momentum-space Hamiltonian, while the $k_{y}$-direction corresponds to the $\Gamma$-$M$ direction. The parameters $v_{j}=\sqrt{3}a\gamma_{j}/2$ are given in terms of the graphene lattice constant, $a=2.46$\mathrm{\SIUnitSymbolAngstrom}$$, and the hoppings, $\gamma_{j}$. Moreover, $\delta$ is a sublattice potential and $D$ is an external displacement field. For our simulations, we use $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\delta)=(3.1,0.38,-0.015,-0.29,-0.141,-0.0105)\,\text{eV}$ [52].
Diagonalizing $H_{0}$ in Eq. (1) shows that the low-energy conduction and valence bands in RTLG originate primarily from the weakly-coupled $A_{1}$ and $B_{4}$ sublattices. These sublattices of the top and bottom layers are only weakly coupled by interlayer hopping, resulting in low-energy bands that touch near the $K$ or $K^{\prime}$ valley. The application of a displacement field, $\Delta_{1}$, opens an energy gap and flattens these lowest bands, leading to a significant enhancement of the density of states, see Fig. 1(a) and (b). Furthermore, a finite $\Delta_{1}$ reverses the band curvature near $\boldsymbol{k}=0$. As a result, the Fermi surface can display multiple connected components depending on $\Delta_{1}$ and the electron density $n_{e}$, see Fig. 1(c).
@@
$\begin{split}H_{\text{eff}}&=\sum_{\boldsymbol{k}}\varepsilon_{\boldsymbol{k}}\,c^{\dagger}_{\boldsymbol{k}}c_{\boldsymbol{k}}\\ &+\frac{1}{2\Omega}\sum_{\boldsymbol{k},\boldsymbol{k}^{\prime},\boldsymbol{q}}\tilde{V}_{\boldsymbol{q}}\,F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}\,F_{\boldsymbol{k}^{\prime},\boldsymbol{k}^{\prime}-\boldsymbol{q}}\;c_{\boldsymbol{k}}^{\dagger}c_{\boldsymbol{k}+\boldsymbol{q}}c_{\boldsymbol{k}^{\prime}}^{\dagger}c_{\boldsymbol{k}^{\prime}-\boldsymbol{q}}\end{split}$ (2)
@@
The first term describes the dispersion, $\varepsilon_{\boldsymbol{k}}$, of the lowest conduction band obtained from $H_{0}$ with $c_{\boldsymbol{k}}$ denoting the electron annihilation operator of that band.
@@
Figure 2: (a) Magnitude of the superconducting order parameter, $\Delta_{\boldsymbol{k},\boldsymbol{q}}$, for the Fermi surface shown as an inset in the zero-temperature limit. (b) Phase of the superconducting order parameter, showing a $p+ip$ winding. (c) Quasiparticle dispersion, $E_{\boldsymbol{k,\boldsymbol{q}}}$, associated with the superconducting order parameter shown in (a) and (b). (d) The quasiparticle dispersion is asymmetric under $k_{x}\rightarrow-k_{x}$. Hence, RTLG realizes a nonreciprocal superconductor.
@@
The second term captures interaction effects by projecting a repulsive Coulomb interaction onto the lowest conduction band and accounting for screening effects at the level of the random phase approximation (RPA). This projection introduces the structure factors, $F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}=\langle\psi_{\boldsymbol{k}}|\psi_{\boldsymbol{k}+\boldsymbol{q}}\rangle$, that describe the overlap of Bloch states in the lowest conduction band. Moreover, the RPA-screened Coulomb potential is,
$\tilde{V}_{\boldsymbol{q}}=\frac{V_{\boldsymbol{q}}}{1-\Pi_{\mathbf{q}}V_{\boldsymbol{q}}}.$ (3)
@@
where $V_{\boldsymbol{q}}=\frac{e^{2}}{2\epsilon}\frac{\tanh(|\boldsymbol{q}|d)}{|\boldsymbol{q}|}$ denotes the gate-screened Coulomb potential with $d$ being the distance to the metallic gates and $\epsilon$ the dielectric permittivity of the hexagonal boron nitride encapsulation. We take $d=20\,\text{nm}$ and $\epsilon=6\epsilon_{0}$ with $\epsilon_{0}$ the vacuum permittivity. The polarization function, $\Pi_{\boldsymbol{q}}$, entering Eq. (3) is given by, $\Pi_{\boldsymbol{q}}=\frac{1}{\Omega}\sum_{\boldsymbol{k}}\frac{n_{F}(\varepsilon_{\boldsymbol{k}+\boldsymbol{q}})-n_{F}(\varepsilon_{\boldsymbol{k}})}{\varepsilon_{\boldsymbol{k}+\boldsymbol{q}}-\varepsilon_{\boldsymbol{k}}}|F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}|^{2}$, where $n_{F}$ is the Fermi-Dirac distribution and $\Omega$ the area of the system.
We will now use this effective Hamiltonian to explain the emergence of nonreciprocal superconductivity and an intrinsic SDE in RTLG.
@@
Nonreciprocal superconductivity. We begin by discussing the emergence of nonreciprocal superconductivity. Nonreciprocal superconductors are a recently proposed class of superconductors in which the breaking of time-reversal and inversion symmetry leads to an asymmetric quasiparticle dispersion [53]. A possible mechanism for such nonreciprocity is finite-momentum pairing. However, as noted in previous works, the Cooper-pair momentum $\boldsymbol{q}$ (defined relative to the $K$ or $K^{\prime}$ valley) in RTLG can be small [44, 45]. We will now explicitly show that the primary origin of the asymmetric quasiparticle dispersion in RTLG is the intrinsic asymmetry of the normal-state band structure, $\varepsilon_{\boldsymbol{k}}\neq\varepsilon_{-\boldsymbol{k}}$. Nevertheless, a finite Cooper-pair momentum can still qualitatively modify the symmetry properties of the quasiparticle dispersion.
@@
To describe the superconductivity in RTLG, we consider, inspired by earlier works [42, 43, 44, 45, 46, 47], a Kohn-Luttinger mechanisms [54, 55, 51, 56, 57]. Specifically, while the bare Coulomb potential, $V_{\boldsymbol{q}}$, is nominally repulsive, the screening effects described by Eq. (3) can lead to an “overscreening” regime, where the effective interaction becomes attractive. This effective attraction can be verified by transforming $\tilde{V}_{\mathbf{q}}$ to real space.
@@
We will assume that this effective attraction can lead to pairing between momentum states $\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$ and $-\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$ where $\boldsymbol{q}$ is the Cooper-pair momentum. While we retain the possibility of a finite $\boldsymbol{q}$, we emphasize that it is not a necessary ingredient for the nonreciprocal superconductivity and SDE. We then perform a mean-field decoupling of the interaction in Eq. (2) in the Cooper channel. The superconducting order parameter is determined self-consistently as,
@@
$\begin{split}\Delta_{\boldsymbol{k},\boldsymbol{q}}=&-\sum_{\boldsymbol{k}^{\prime}}g_{\boldsymbol{k},\boldsymbol{k}^{\prime},\boldsymbol{q}}\,\frac{\Delta_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2\tilde{E}_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}\\ \times&\frac{1}{2}\left\{\tanh\left(\frac{\beta E_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2}\right)+\tanh\left(\frac{\beta E_{-\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2}\right)\right\}.\end{split}$ (4)
@@
Next, we solve for $\Delta_{\boldsymbol{k},\boldsymbol{q}}$ self-consistently, imposing a fixed electron density $n_{e}=\sum_{\boldsymbol{k}}\langle c^{\dagger}_{\boldsymbol{k}}c_{\boldsymbol{k}}\rangle/\Omega$ and varying $\boldsymbol{q}$. The realized superconducting order, $\Delta_{\boldsymbol{k},\boldsymbol{q}}$, is determined by minimizing the free energy,
@@
Figure 3: (a) Free energy, $F$, versus $\boldsymbol{q}=(q_{x},q_{y})$ for the annulus Fermi surface shown in Fig. 2. Here, $F_{\text{min}}\equiv\text{min}_{\boldsymbol{q}}\{F(\boldsymbol{q})\}.$ (b) Corresponding supercurrent magnitude, $I$, as a function of $\boldsymbol{q}$. Here, $I_{\text{max}}\equiv\text{max}_{\boldsymbol{q}}\{I(\boldsymbol{q})\}$. (c) Critical currents in the forward direction ($I_{c,+}$) and reverse direction ($I_{c,-}$) for current direction given by an angle $\theta$ relative to the $x$-direction. (d) Diode efficiency for a current direction given by an angle $\theta$.
@@
First, we assume $\boldsymbol{q}=0$. The normal state spectrum is invariant under two three-fold rotations ($2C_{3}$) and three two-fold rotations ($3C_{2}^{\prime}$) around the in-plane axis pointing along $\theta\in\{0^{\circ},60^{\circ},120^{\circ}\}$. Consider a $\boldsymbol{k}$ pointing along one of the directions $\theta\in\{90^{\circ},150^{\circ},30^{\circ}\}$. Then, $C_{2}^{\prime}\boldsymbol{k}=-\boldsymbol{k}$ and hence $\varepsilon_{-\boldsymbol{k}}=\varepsilon_{C_{2}^{\prime}\boldsymbol{k}}=\varepsilon_{\boldsymbol{k}}$. As a result, $\xi_{\boldsymbol{k},\boldsymbol{q},-}=0$ and $E_{\boldsymbol{k},\boldsymbol{q}}=E_{-\boldsymbol{k},\boldsymbol{Q}}$ (since $|\Delta_{\boldsymbol{k},\boldsymbol{q}}|=|\Delta_{-\boldsymbol{k},\boldsymbol{q}}|$). Consequently, if $\boldsymbol{q}=0$, the quasiparticle spectrum is symmetric along the directions $\theta\in\{90^{\circ},150^{\circ},30^{\circ}\}$, but asymmetric otherwise.
@@
Now assume $\boldsymbol{q}\neq 0$. In this case, a vanishing asymmetry of the quasiparticle spectrum requires both $C_{2}^{\prime}\boldsymbol{k}=-\boldsymbol{k}$ and $C_{2}^{\prime}\boldsymbol{q}=\boldsymbol{q}$. In our simulations, we found that $\boldsymbol{q}$ typically pointed along the $\theta=0^{\circ}$ directions so that the two conditions can only be satisfied for momenta $\boldsymbol{k}$ with $\theta=90^{\circ}$. Consequently, the quasiparticle dispersion can be strictly symmetric only along $\theta=90^{\circ}$; in all other directions, it remains asymmetric.

### v2

@@ Abstract
Recent experiments have revealed that superconductivity in rhombohedral tetralayer graphene can emerge from a valley-polarized and, hence, chiral normal state. The interplay of pairing and the reduced normal-state symmetries sparked widespread interest. In this work, we demonstrate within a microscopic theoretical formalism that this stabilizes a non-reciprocal superconducting state. Our results are based on a fully self-consistent framework for determining the superconducting order parameter from a Kohn-Luttinger mechanism. We show that the system displays a sizeable intrinsic superconducting diode effect, i.e., without the need for applying external magnetic fields, which is also highly tunable by an external displacement field. Moreover, we find that the angular dependence of the critical current is highly sensitive to the Fermi surfaces of the normal state. Hence, our results suggest that the critical current could provide insights into the type of Fermi surface topology from which superconductivity arises.
@@
Non-reciprocal phenomena in superconducting systems that break time-reversal and inversion symmetry have recently attracted a great deal of attention [1]. A key example is the superconducting diode effect (SDE), characterized by a critical current that depends on the current-bias direction [2, 3, 4, 5, 6, 7, 8]. One route to achieving such a SDE relies on applying an external magnetic field [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] or flux [19, 20, 21, 22, 23, 24, 25, 26] to break time-reversal symmetry. However, it is also possible for materials to break the relevant symmetries spontaneously at the microscopic level, giving rise to an intrinsic SDE [27, 28, 29, 30, 31, 32, 33]. Such an intrinsic SDE could enable field-free non-reciprocal circuit elements [34, 35, 36, 37, 38, 39] and serve as a probe for symmetry-breaking in superconducting materials.
@@
In an exciting recent development, superconductivity has been discovered in rhombohedral tetralayer graphene (RTLG) [40, 41]. A unique aspect of RTLG is that a superconducting phase can emerge from a spin- and valley-polarized “quarter-metal” parent state and persists under large in-plane magnetic fields [40]. These observations point to a possible spin-polarized pairing with a chiral $p+ip$ order parameter, as suggested by theoretical works [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Importantly, the valley polarization in the normal state breaks time-reversal symmetry and in-plane inversion (i.e., two-fold rotational symmetry) [28], making RTLG a promising platform for non-reciprocal superconducting phenomena. Indeed, an observation of an intrinsic SDE in RTLG has recently been reported [40]. However, a proper and quantitative understanding of the diode effect beyond symmetry arguments requires an explicit microscopic calculation.
@@
Figure 1: (a) Dispersion of RTLG along the $k_{x}$-direction for different displacement fields, $\Delta_{1}$, showing a band flattening around $k_{x}=0$. (b) Density of states (DOS) versus $\Delta_{1}$ and the carrier density, $n_{e}$ (relative to charge neutrality), showing a significant enhancement due to the band flattening. (c) Fermi surfaces for ($\Delta_{1},n_{e}$), from left to right, set to $(72\,\text{meV},0.574)$, $(67.5\,\text{meV},0.386)$, $(65\,\text{meV},0.414)$, $(63\,\text{meV},0.254)$.
@@
Here, we fill this gap by theoretically demonstrating and studying the emergence of nonreciprocal superconductivity and an intrinsic SDE in RTLG. Our results are based on a fully self-consistent framework for determining the superconducting order parameter from a Kohn-Luttinger mechanism to superconductivity and show that large diode efficiencies are achievable and highly tunable by a displacement field. Moreover, we find that both the diodicity and the critical current in general show a characteristic angular dependence with enhanced lobes, which depend on the Fermi surface topology of the normal state. We further show that the same intrinsic non-reciprocity produces an asymmetric current–voltage characteristic in a metal–superconductor junction. Taken together, our findings suggest that measurements of supercurrent and conductance asymmetries could provide insights into the type of Fermi surface topology from which superconductivity arises.
@@
Model. To describe spin- and valley-polarized RTLG, we use an $8$-band $\mathbf{k}\cdot\mathbf{p}$ model for the electronic states near the $K$ or $K^{\prime}$ valley (labeled by $\tau=\pm$). The Bloch Hamiltonian, $h_{0}$, capturing the non-interacting bands, written in the basis $(A_{1},B_{1},A_{2},B_{2},A_{3},B_{3},A_{4},B_{4})$, where $A_{\ell}$ and $B_{\ell}$ denote the sublattices of the $\ell^{\text{th}}$ layer, is given by [53, 54, 55],
@@
$\begin{pmatrix}-\Delta_{1}+\delta&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi&0&\frac{\gamma_{2}}{2}&0&0\\ v_{0}\pi&-\Delta_{1}&\gamma_{1}&v_{4}\pi^{\dagger}&0&0&0&0\\ v_{4}\pi&\gamma_{1}&-\frac{\Delta_{1}}{3}&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi&0&\frac{\gamma_{2}}{2}\\ v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&-\frac{\Delta_{1}}{3}&\gamma_{1}&v_{4}\pi^{\dagger}&0&0\\ 0&0&v_{4}\pi&\gamma_{1}&\frac{\Delta_{1}}{3}&v_{0}\pi^{\dagger}&v_{4}\pi^{\dagger}&v_{3}\pi\\ \frac{\gamma_{2}}{2}&0&v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&\frac{\Delta_{1}}{3}&\gamma_{1}&v_{4}\pi^{\dagger}\\ 0&0&0&0&v_{4}\pi&\gamma_{1}&\Delta_{1}&v_{0}\pi^{\dagger}\\ 0&0&\frac{\gamma_{2}}{2}&0&v_{3}\pi^{\dagger}&v_{4}\pi&v_{0}\pi&\Delta_{1}+\delta\end{pmatrix}.$
Here $\pi=\tau k_{x}+ik_{y}$ and the momentum $\boldsymbol{k}=(k_{x},k_{y})$ is measured relative to $K$ or $K^{\prime}$ valley. The $k_{x}$-direction correspond to the $\Gamma$-$K$ direction of the full momentum-space Hamiltonian, while the $k_{y}$-direction corresponds to the $\Gamma$-$M$ direction. The parameters $v_{j}=\sqrt{3}a\gamma_{j}/2$ are given in terms of the graphene lattice constant, $a=2.46$\mathrm{\SIUnitSymbolAngstrom}$$, and the hoppings, $\gamma_{j}$. Moreover, $\delta$ is a sublattice potential and $\Delta_{1}$ is an external displacement field. For our simulations, we use $(\gamma_{0},\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4},\delta)=(3.1,0.38,-0.015,-0.29,-0.141,-0.0105)\,\text{eV}$ [56].
Diagonalizing $h_{0}$ in Eq. (1) shows that the low-energy conduction and valence bands in RTLG touch near the $K$ or $K^{\prime}$ valley and originate primarily from the weakly-coupled $A_{1}$ and $B_{4}$ sublattices. The application of a displacement field, $\Delta_{1}$, opens an energy gap and flattens these lowest bands, leading to a significant enhancement of the DOS, see Fig. 1(a) and (b). Furthermore, a finite $\Delta_{1}$ reverses the band curvature near $\boldsymbol{k}=0$. As a result, the Fermi surface can display multiple connected components depending on $\Delta_{1}$ and the electron density $n_{e}$, see Fig. 1(c).
@@
$\begin{split}H_{\text{eff}}&=\sum_{\boldsymbol{k}}\varepsilon_{\boldsymbol{k}}\,c^{\dagger}_{\boldsymbol{k}}c^{\phantom{\dagger}}_{\boldsymbol{k}}\\ &+\frac{1}{2\Omega}\sum_{\boldsymbol{k},\boldsymbol{k}^{\prime},\boldsymbol{q}}\tilde{V}_{\boldsymbol{q}}\,F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}\,F_{\boldsymbol{k}^{\prime},\boldsymbol{k}^{\prime}-\boldsymbol{q}}\;c_{\boldsymbol{k}}^{\dagger}c^{\phantom{\dagger}}_{\boldsymbol{k}+\boldsymbol{q}}c_{\boldsymbol{k}^{\prime}}^{\dagger}c^{\phantom{\dagger}}_{\boldsymbol{k}^{\prime}-\boldsymbol{q}}\end{split}$ (2)
@@
The first term describes the dispersion, $\varepsilon_{\boldsymbol{k}}$, of the lowest conduction band obtained from $h_{0}$ with $c_{\boldsymbol{k}}$ denoting the electron annihilation operator of the active spin and valley flavor in that band.
@@
Figure 2: (a) Magnitude of the superconducting order parameter, $\Delta_{\boldsymbol{k},\boldsymbol{q}}$, for the Fermi surface shown as an inset in the zero-temperature limit. (b) Phase of the superconducting order parameter, showing a $p+ip$ winding. Here, we have chosen a gauge such that the wavefunction of the lowest conduction band has real-valued inner product with the wavefunction at $\boldsymbol{k}=0$, i.e., $\innerproduct{\psi_{\boldsymbol{k}}}{\psi_{\boldsymbol{0}}}$ is real. (c) Quasiparticle dispersion, $E_{\boldsymbol{k,\boldsymbol{q}}}$, associated with the superconducting order parameter shown in (a) and (b). (d) The quasiparticle dispersion is asymmetric under $k_{x}\rightarrow-k_{x}$. Hence, RTLG realizes a nonreciprocal superconductor.
@@
The second term captures interaction effects by projecting a repulsive Coulomb interaction onto the lowest conduction band and accounting for screening effects via a random phase approximation (RPA). This projection introduces the form factors, $F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}=\langle\psi_{\boldsymbol{k}}|\psi_{\boldsymbol{k}+\boldsymbol{q}}\rangle$, that describe the overlap of Bloch states in the lowest conduction band. The screened Coulomb potential is
$\tilde{V}_{\boldsymbol{q}}=\frac{V_{\boldsymbol{q}}}{1-\Pi_{\mathbf{q}}V_{\boldsymbol{q}}},$ (3)
@@
where $V_{\boldsymbol{q}}=\frac{e^{2}}{2\epsilon}\frac{\tanh(|\boldsymbol{q}|d)}{|\boldsymbol{q}|}$ denotes the gate-screened Coulomb potential with $d$ being the distance to the metallic gates and $\epsilon$ the dielectric permittivity of the hexagonal boron nitride encapsulation. We take $d=20\,\text{nm}$ and $\epsilon=6\epsilon_{0}$ with $\epsilon_{0}$ the vacuum permittivity. The polarization function, $\Pi_{\boldsymbol{q}}$, entering Eq. (3) is given by $\Pi_{\boldsymbol{q}}=\frac{1}{\Omega}\sum_{\boldsymbol{k}}\frac{n_{F}(\varepsilon_{\boldsymbol{k}+\boldsymbol{q}})-n_{F}(\varepsilon_{\boldsymbol{k}})}{\varepsilon_{\boldsymbol{k}+\boldsymbol{q}}-\varepsilon_{\boldsymbol{k}}}|F_{\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q}}|^{2}$, where $n_{F}$ is the Fermi-Dirac distribution and $\Omega$ the area of the system. We will now use this effective Hamiltonian to explain the emergence of superconductivity, study the resulting order parameter, and demonstrate its non-reciprocal signatures, including a sizable intrinsic SDE.
@@
Nonreciprocal superconductivity. We begin by discussing nonreciprocal superconductivity. Nonreciprocal superconductors are a recently proposed class of superconductors in which the breaking of time-reversal and inversion symmetry leads to an asymmetric quasiparticle dispersion [57]. Here, we will explicitly show that in RTLG this asymmetric quasiparticle dispersion arises from the asymmetry of the normal-state band structure, $\varepsilon_{\boldsymbol{k}}\neq\varepsilon_{-\boldsymbol{k}}$, and discuss how a finite Cooper-pair momentum (relative to $K$ or $K^{\prime}$[44, 45] can influence the the symmetry properties of quasiparticle dispersion.
@@
To describe the superconductivity in RTLG, we consider, inspired by earlier works [42, 43, 44, 45, 46, 47], a Kohn-Luttinger mechanisms [58, 59, 55, 60, 61]. Specifically, while the bare Coulomb potential, $V_{\boldsymbol{q}}$, is nominally repulsive, the screening effects described by Eq. (3) can lead to an “overscreening” regime, where the effective interaction becomes attractive in some channels. This effective attraction can be verified by transforming $\tilde{V}_{\mathbf{q}}$ to real space.
@@
We will assume that this effective attraction can lead to pairing between momentum states $\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$ and $-\boldsymbol{k}+\frac{\boldsymbol{q}}{2}$ where $\boldsymbol{q}$ is the Cooper-pair momentum. While we retain the possibility of a finite $\boldsymbol{q}$, we emphasize that it is not a necessary ingredient for the nonreciprocal superconductivity and SDE. We then perform a mean-field decoupling of Eq. (2) in the Cooper channel. The superconducting order parameter is determined self-consistently as,
@@
$\begin{split}\Delta_{\boldsymbol{k},\boldsymbol{q}}=&-\frac{1}{\Omega}\sum_{\boldsymbol{k}^{\prime}}g_{\boldsymbol{k},\boldsymbol{k}^{\prime},\boldsymbol{q}}\,\frac{\Delta_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2\tilde{E}_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}\\ \times&\frac{1}{2}\left\{\tanh\left(\frac{\beta E_{\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2}\right)+\tanh\left(\frac{\beta E_{-\boldsymbol{k}^{\prime},\boldsymbol{q}}}{2}\right)\right\}.\end{split}$ (4)
@@
Figure 3: (a) Free energy, $F$, versus $\boldsymbol{q}=(q_{x},q_{y})$ for the annulus Fermi surface shown in Fig. 2. Here, $F_{\text{min}}\equiv\text{min}_{\boldsymbol{q}}\{F(\boldsymbol{q})\}.$ (b) Corresponding supercurrent magnitude, $I$, versus $\boldsymbol{q}$. Here, $I_{\text{max}}\equiv\text{max}_{\boldsymbol{q}}\{I(\boldsymbol{q})\}$. (c) Critical currents in the forward direction ($I^{+}_{c}$) and reverse direction ($I^{-}_{c}$) for current direction given by an angle $\theta$ relative to the $x$-direction. (d) Diode efficiency for a current direction given by an angle $\theta$.
Next, we solve for $\Delta_{\boldsymbol{k},\boldsymbol{q}}$ self-consistently, imposing a fixed electron density $n_{e}=\sum_{\boldsymbol{k}}\langle c^{\dagger}_{\boldsymbol{k}}c_{\boldsymbol{k}}\rangle/\Omega$ and varying $\boldsymbol{q}$. The realized superconducting order parameter, $\Delta_{\boldsymbol{k},\boldsymbol{q}}$, is determined by minimizing the free energy,
@@
First, assume $\boldsymbol{q}=0$. On top of the aforementioned $C_{3}$ symmetry, the normal state spectrum is invariant under reflection $\sigma_{v}$ at the mirror planes along $\theta\in\{0^{\circ},60^{\circ},120^{\circ}\}$.
@@
Consider a $\boldsymbol{k}$ pointing along one of the perpendicular directions $\theta\in\{90^{\circ},150^{\circ},30^{\circ}\}$. Then, $\sigma_{v}\boldsymbol{k}=-\boldsymbol{k}$ and hence $\varepsilon_{-\boldsymbol{k}}=\varepsilon_{\sigma_{v}\boldsymbol{k}}=\varepsilon_{\boldsymbol{k}}$. As a result, $\xi_{\boldsymbol{k},\boldsymbol{q},-}=0$ and $E_{\boldsymbol{k},\boldsymbol{q}}=E_{-\boldsymbol{k},\boldsymbol{q}}$ (since $|\Delta_{\boldsymbol{k},\boldsymbol{q}}|=|\Delta_{-\boldsymbol{k},\boldsymbol{q}}|$). Thus, the quasiparticle spectrum is symmetric along $\theta\in\{90^{\circ},150^{\circ},30^{\circ}\}$, but asymmetric otherwise.
@@
Now assume $\boldsymbol{q}\neq 0$. In this case, a vanishing asymmetry of the quasiparticle spectrum requires both $\sigma_{v}\boldsymbol{k}=-\boldsymbol{k}$ and $\sigma_{v}\boldsymbol{q}=\boldsymbol{q}$. For $\boldsymbol{q}=(q_{x},0)^{T}$, i.e., pointing along $\theta=0^{\circ}$, the two conditions can only be satisfied for momenta $\boldsymbol{k}$ with $\theta=90^{\circ}$. Consequently, the quasiparticle dispersion can be strictly symmetric only along $\theta=90^{\circ}$; in all other directions, it remains asymmetric. Note, however, that as a result of the numerically small value of $q_{x}$, the dispersion shown in Fig. 2(c) still looks approximately symmetric along $\theta=90^{\circ},150^{\circ}$, and $30^{\circ}$.